# Estimating the linear relation between variables that are never jointly observed

### v3

@@ Abstract
In modern experimental science there is a commonly encountered problem of estimating the coefficients of a linear regression in the context where the variables of interest can never be observed simultaneously. Assuming that the global experiment can be decomposed into sub-experiments with distinct first moments, we propose two estimators of the linear regression that take this additional information into account. We consider an estimator based on moments, and an estimator based on optimal transport theory. These estimators are proven to be consistent as well as asymptotically Gaussian under weak hypotheses. The asymptotic variance has no explicit expression, except in some particular cases, for which reason a stratified bootstrap approach is developed to build confidence intervals for the estimated parameters, whose consistency is also shown. A simulation study, assessing and comparing the finite sample performances of these estimators, demonstrated the advantages of the bootstrap approach in multiple realistic scenarios. An application to in vivo experiments, conducted in the context of studying radio-induced adverse effects on mice, revealed important relationships between the biomarkers of interest that could not be identified with the considered naive approach.
@@ 1 Introduction
In order to study the effects of a certain treatment on a living organism, in vivo experiments are often conducted. In the context of a complex organism response, scientists may be interested in studying multiple variables describing the effect on different scales. In particular, such variables of interest often include a macroscopic biomarker, only available through in vivo data, and a microscopic biomarker that can also be observed on a cellular level (i.e. in vitro). The interest in this case lies in predicting the former with the latter. For instance, in the context of studying the adverse effects induced by radiotherapy on healthy tissues, the potential outcomes of interest manifest in the form of the severity of macroscopic lesions, and predictors such as gene expression. These measures often require sacrificing the animals that they are taken from. As a result, the quantities of interest cannot be observed on the same animals. Since the goal is to establish relationships between these variables, a problem of statistical data fusion arises.
@@
Figure 1.1: Schematic representation of the design of an in vivo experiment studying the effect of irradiated volume.
@@
An illustrative example of an in vivo experiment where the variables of interest are never observed simultaneously is presented in Figure 1.1. In this experiment, mice are irradiated on the lungs with different volume, with a goal of studying the role of irradiated volume in the appearance of radio-induced adverse effects. The latter are assessed by measuring septal thickening, a histological marker of lung injuries. The other variable that is measured with the purpose of predicting the adverse effect-related variable is the expression of multiple pro-inflammatory genes. As shown in Figure 1.1, there are two independent cohorts in the study, one is used to measure gene expression, whereas the other for measuring septal thickness. This is a results of both measures being of destructive nature, which does not allow them to be taken on the same animals.
@@
Comparing distributions of the measurements arising from the two cohorts, such as those presented in Figure 1.2, one may suspect a correlation or even a linear relationship between the variables. In order to assert whether such a relationship exists, one has to connect two variables that are never observed simultaneously, which translates into solving a data fusion problem. In this example, there are four groups indicating time points (1, 3, 6 and 12 months after irradiation) when the corresponding animals are sacrificed and the measurements are taken. Thus, the categorical variable indicating the time point, which is known for every observation, can be used as an additional variable in order to link the predictor and the predicted variables between each other.
@@
Figure 1.2: Distribution of the data, collected from the irradiated patch under SBRT with 3 mm beam size: the expression of the gene IL6 on the left, and septal thickness on the right. The measurements were made 1, 3, 6 and 12 months after irradiation.
@@
The task of linking variables that are never jointly observed cannot be approached as a typical missing values problem, since most methods for inference on incomplete data require a sufficient overlap, which is completely absent in the case treated here. As a result, any approaches applying frameworks such as multiple imputation in the context of data fusion are unsuitable for our application. For example, Carrig et al., (2015) use multiple imputation to integrate disparate datasets allowing for the absence of the overlap, but requiring a calibration dataset, where all the variables of interest must be jointly observed.
@@
Other approaches to data fusion available in literature include factor analysis (Cudeck,, 2000), statistical matching (Mitsuhiro and Hoshino,, 2020), Bayesian network inference (Triantafillou et al.,, 2010; Tsamardinos et al.,, 2012) and Gaussian Markov combinations (Massa and Riccomagno,, 2017). These methods are designed for linking variables that are not observed simultaneously through covariates, present for both variables of interest. This corresponds to the properties of the in vivo data described above. However, the covariates in these approaches are random variables, typically continuous, and often assumed to be Gaussian, which is the case in Cudeck, (2000) and Massa and Riccomagno, (2017). The grouping variable available through in vivo experiments cannot be represented in a continuous form since categories such as control and sham make it impossible to assume continuity and normality. The Bayesian network approaches introduced by Triantafillou et al., (2010) and Tsamardinos et al., (2012), aimed at inferring binary causal relationships between variables, are more suitable for large datasets with a high number of covariates. Finally, currently available research in statistical matching addresses such aspects as not-at-random missingness (Mitsuhiro and Hoshino,, 2021) and high dimensionality (Mitsuhiro and Hoshino,, 2020). This approach is based on the idea of comparing distances between the covariates from the datasets of interest, which cannot be done by taking the group variable as the covariate. It can be noted that the goal of the aforementioned examples in statistical matching is to group individuals before imputation, which is not necessary in our case since the groups are already known.
In this work we propose an approach to estimate the relationship between the variables that are not simultaneously observed under the experimental setting described above, based on a conditional model assuming linear relationship within every component, defined by the grouping variable that is assumed to be present. An estimator derived with the method of moments as well as optimal transport solution using Wasserstein distance are considered for the real data problem in question. Both approaches do not require any overlap between the two cohorts of the experiment, and are based on weak assumptions ensuring model identifiability and moments existence. To construct confidence intervals, we propose both asymptotic and bootstrap-based results, the former being often unfeasible in practice, with the latter demonstrating practical advantages in numerous realistic settings.
@@ 2 Identification approaches
We consider a real random variable $Y$ and a vector of $d$ real valued random regressors $\mathbf{X}=(X_{1},\ldots,X_{d})$ and suppose that the following linear regression hold
$Y$ $=\beta_{0}+\sum_{j=1}^{d}\beta_{j}X_{j}+\epsilon.$ (2.1)
@@
In the absence of additional information and without any strong additional hypothesis, the parameters $(\beta_{0},\beta_{1},\ldots,\beta_{d})$ and the variance of the noise $\sigma^{2}_{\epsilon}$ cannot be identified. Indeed, if for example $X_{1}$ is centered with symmetric distribution, the coefficient $\beta_{1}$ can only be determined up to sign change since $\beta_{1}X_{1}$ and $\beta_{1}(-X_{1})$ have the same distribution.
To deal with this identification issue, we consider that we can perform different experiments in which the mean value of $X$ is allowed to vary. For that, we suppose that there are $K$ groups (corresponding to $K$ different experiments), defined by a discrete variable $G$ taking values in $\{1,\cdots,K\}$ observed simultaneously with $Y$ and with $X$. This means that we have now access to $(\mathbf{X},G)$ and $(Y,G)$ but not to $(\mathbf{X},Y,G)$. We also suppose that $\epsilon$ is independent of $G$.
Given $G=k$, for $k=1,\ldots,K$, we denote by $\mu_{Y}^{k}={\mathbb{E}}(Y|G=k)$ and $\mu_{X_{j}}^{k}={\mathbb{E}}(X_{j}|G=k)$, $j=1,\ldots,d$, the expected values within each group.
We now present two different approaches developed to identify the vector $\boldsymbol{\beta}=(\beta_{0},\ldots,\beta_{d})$ of unknown regression coefficients and the noise variance $\sigma_{\epsilon}^{2}$, taking account of the additional information given by the discrete variable $G$.
@@ 2.1 Moment approach
A first and simple approach is based on first moments identification. Taking the conditional expectation, given $G=k$, in (2.1), we have for $k=1,\ldots,K$,
$\mu_{Y}^{k}$ $=\beta_{0}+\sum_{j=1}^{d}\beta_{j}\mu_{X_{j}}^{k},$ (2.2)
since the residual term $\epsilon$ is supposed to satisfy $\mathbb{E}(\epsilon|G=k)=0$ for $k=1,\ldots,K$.
We denote by $\boldsymbol{\mu}_{1,X}$ the $K\times(d+1)$ design matrix, whose $k$th row is equal to $(1,\boldsymbol{\mu}_{X}^{k\top})$ with $\boldsymbol{\mu}_{X}^{k}=(\mu_{X_{1}}^{k},\cdots,\mu_{X_{d}}^{k})^{\top}$, by $\boldsymbol{\mu}_{Y}$ the $K$ dimensional vector with elements $(\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$. The $K$ linear equations in (2.2) can we written in a matrix way, $\boldsymbol{\mu}_{Y}=\boldsymbol{\mu}_{1,X}\boldsymbol{\beta}$.
We introduce the following assumption, guaranteeing the identifiability of the model parameters:
@@
Lemma 2.1.
If model (2.1) holds and assumption $\mathbf{H}_{1}$ is fulfilled, $\boldsymbol{\beta}$ is uniquely identified in terms of the conditional first order moments of $\mathbf{X}$ and $Y$ given $G$,
@@ 2.2 Optimal transport approach
The second approach is based on optimal transport, in particular on the idea of estimating the linear transformation of the distribution of $\mathbf{X}$ that is the closest to that of $Y$ with respect to the Wasserstein distance (see Panaretos and Zemel, (2019) for a general introduction for statisticians).
The optimal transport map T between Gaussian measures on $\mathbb{R}^{d}$ is linear, and the Wasserstein distance of order 2 between two Gaussian distributions $D_{1}$ and $D_{2}$, with $D_{1}\sim\mathcal{N}(\boldsymbol{\mu}_{1},\boldsymbol{\Gamma}_{1})$ and $D_{2}\sim\mathcal{N}(\boldsymbol{\mu}_{2},\boldsymbol{\Gamma}_{2})$, is equal to
@@
If we assume that, given $G=k$, $\mathbf{X}$ is a Gaussian random vector and $\epsilon$ is Gaussian, we have, if the linear model (2.1) holds, that $Y$ is also Gaussian given $G=k$, with expectation $\mu_{Y}^{k}=\beta_{0}+\sum_{j=1}^{d}\beta_{j}\mu_{X_{j}}^{k}$ and variance $\sigma_{\epsilon}^{2}+\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}$, where $\boldsymbol{\Gamma}_{X}^{k}$ is the variance matrix of $\mathbf{X}$ given $G=k$. Thus, the Wasserstein distance between $D_{\gamma}$, the distribution of $\gamma_{0}+\boldsymbol{\gamma}_{-0}^{\top}\mathbf{X}+\epsilon$, and $D_{Y}$, the distribution of $Y$, is equal to
@@
$\varphi(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\mu_{Y}^{k}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k})^{2}+\left(\sigma_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (2.3)
which evaluates the weighted Wasserstein distance between $Y$ and a linear combination of the $\mathbf{X}$ variables, contaminated by a Gaussian noise, with variance $\sigma^{2}$. We state, without proof, the following Lemma ensuring that the parameters $\boldsymbol{\beta}$ and $\sigma_{\epsilon}^{2}$ can be identified under general conditions.
Lemma 2.2.
If model (2.1) holds and if assumption $\mathbf{H}_{1}$ is fulfilled, $\varphi(\boldsymbol{\gamma},\sigma^{2})$ has its unique minimum at $\boldsymbol{\gamma}=\boldsymbol{\beta}$ and $\sigma^{2}=\sigma_{\epsilon}^{2}$.
@@ 3 Sampled data and estimators
We suppose that experiments are made for $K\geq 2$ different groups and that for each group $k$, for $k=1,\ldots,K$, we have two independent samples $(Y^{k}_{1},\ldots,Y^{k}_{n_{y}^{k}})$ and $(X^{k}_{j,1},\ldots,X^{k}_{j,n_{x}^{k}})_{j=1,\ldots,d}$, with sizes $n_{y}^{k}$ and $n_{x}^{k}$. For each unit $i=1,\ldots,n_{x}^{k}$ from group $k$, the vector of covariates is denoted by $\mathbf{X}_{i}^{k}=(X_{1,i},\ldots,X_{d,i})$. We also define $N_{x}=\sum_{k=1}^{K}n_{x}^{k}$ and $N_{y}=\sum_{k=1}^{K}n_{y}^{k}$, the total number of observations of the response $Y$ and the covariates $X_{1},\ldots,X_{d}$.
As shown in Lemma 2.1 and Lemma 2.2, identification of parameter $\boldsymbol{\beta}$ relies on the knowledge of the first two conditional moments, given $G$, of $Y$ and $\mathbf{X}$. For $k=1,\ldots,K$, we denote by $\widehat{\mu}_{Y}^{k}=\frac{1}{n_{y}^{k}}\sum_{i=1}^{n_{y}^{k}}Y^{k}_{i}$ and $\widehat{\mu}_{X_{j}}^{k}=\frac{1}{n_{x}^{k}}\sum_{i=1}^{n_{x}^{k}}X^{k}_{j,i}$ the empirical means within each group $k$ and by $\widehat{\sigma}^{2}_{Y,k}=\frac{1}{n_{y}^{k}}\sum_{i=1}^{n_{y}^{k}}\left(Y_{i}^{k}\right)^{2}-\left(\widehat{\mu}_{Y}^{k}\right)^{2}$, and $\widehat{\boldsymbol{\Gamma}}^{k}_{X}=\frac{1}{n_{x}^{k}}\sum_{i=1}^{n_{x}^{k}}\mathbf{X}_{i}^{k}(\mathbf{X}_{i}^{k})^{\top}-\widehat{\boldsymbol{\mu}}_{X}^{k}(\widehat{\boldsymbol{\mu}}_{X}^{k})^{\top}$, the empirical variances, with $\widehat{\boldsymbol{\mu}}_{X}^{k}=(\widehat{\mu}_{X_{1}}^{k},\ldots,\widehat{\mu}_{X_{d}}^{k})$. We also define the overall empirical mean and variance $\widehat{\mu}_{Y}=\frac{1}{N_{y}}\sum_{k=1}^{K}{n_{y}^{k}}\widehat{\mu}_{Y}^{k}$, $\widehat{\boldsymbol{\mu}}_{X}=\frac{1}{N_{x}}\sum_{k=1}^{K}{n_{x}^{k}}\widehat{\boldsymbol{\mu}}_{X}^{k}$, $\widehat{\sigma}^{2}_{Y}=\frac{1}{N_{y}}\sum_{k=1}^{K}\sum_{i=1}^{n_{y}^{k}}\left(Y_{i}^{k}\right)^{2}-\left(\widehat{\mu}_{Y}\right)^{2}$, $\widehat{\boldsymbol{\Gamma}}_{X}=\frac{1}{N_{x}}\sum_{k=1}^{K}\sum_{i=1}^{n_{x}^{k}}\mathbf{X}_{i}^{k}(\mathbf{X}_{i}^{k})^{\top}-\widehat{\boldsymbol{\mu}}_{X}\widehat{\boldsymbol{\mu}}_{X}^{\top}$.
We denote by $\widehat{\boldsymbol{\mu}}_{1,X}$ the $K\times(d+1)$ matrix, with the first column consisting of ones, and the rest equal to $\widehat{\boldsymbol{\mu}}_{X}$.

### v4

# Bootstrap inference for linear regression between variables that are never jointly observed: application in in vivo experiments

@@ Abstract
In modern experimental science, there is a common problem of estimating the coefficients of a linear regression in a context where the variables of interest cannot be observed simultaneously. When there is a categorical variable that is observed on all statistical units, we consider two estimators of linear regression that take this additional information into account: an estimator based on moments and an estimator based on optimal transport theory. These estimators are shown to be consistent and asymptotically Gaussian under weak hypotheses. The asymptotic variance has no explicit expression, except in some special cases, for which reason a stratified bootstrap approach is developed to construct confidence intervals for the estimated parameters, whose consistency is also shown. A simulation study evaluating and comparing the finite sample performance of these estimators demonstrates the advantages of the bootstrap approach in several realistic scenarios. An application to in vivo experiments, conducted in the context of studying radio-induced adverse effects in mice, revealed important relationships between the biomarkers of interest that could not be identified with the considered naive approach.
@@ 1 Introduction
In vivo experiments are often used to study the effects of a treatment on a living organism. In the context of a complex organism response, scientists may be interested in studying multiple variables that describe the effect at different scales. In particular, such variables of interest often include a macroscopic biomarker that is only available through in vivo data, and a microscopic biomarker that can also be observed at the cellular level (i.e. in vitro). The interest in this case is to use the latter to predict the former. For example, in the context of studying the adverse effects of radiotherapy on healthy tissue, the potential outcomes of interest are the severity of macroscopic lesions and predictors such as gene expression. In preclinical research, these quantities of interest are often observed in different animals from independent cohorts. Since the goal is to establish relationships between these variables, the problem of statistical data fusion arises.
@@
Figure 1: Schematic representation of the design of an in vivo experiment studying the effect of irradiated volume.
@@
An illustrative example of an in vivo experiment where the variables of interest are not observed simultaneously is shown in Figure 1. In this experiment, presented in Bertho et al., (2020), mice are irradiated in the lungs with different volumes to study the role of irradiated volume in the occurrence of radiation-induced adverse effects. The latter are assessed by measuring septal thickening, a histological marker of lung injury. The other variable measured to predict the adverse effect variable is the expression of several pro-inflammatory genes. As shown in Figure 1, there are two independent cohorts in the study, one used to measure gene expression and the other to measure septal thickness.
@@
Comparing the distributions of measurements from the two cohorts, as shown in Figure 2, may suggest a correlation or even a linear relationship between the variables. To establish whether such a relationship exists, it is necessary to link two variables that are not observed on the same statistical units, which is equivalent to solving a data fusion or a file matching problem according to the terminology employed in Little and Rubin, (2002). In this example, there are four groups indicating the time points (1, 3, 6 and 12 months after irradiation) when the measurements were taken on the corresponding animals. Thus, the categorical variable indicating the time point, which is known for each observation, can be used as an additional variable to link the predictor and the predicted variables.
@@
Figure 2: Distribution of the data, collected from the irradiated patch under SBRT with 3 mm beam size: the expression of the gene IL6 on the left, and septal thickness on the right. The measurements were made 1, 3, 6 and 12 months after irradiation.
@@
The task of linking variables that are not observed together cannot be approached as a typical missing value problem, since most methods of inference on incomplete data require sufficient overlap, which is completely absent in the case we are dealing with. As a result, all approaches that use frameworks such as multiple imputation in the context of data fusion are inappropriate for our application. For example, Carrig et al., (2015) use multiple imputation to integrate different datasets, which allows for the absence of overlap, but requires a calibration dataset in which all variables of interest must be jointly observed.
@@
Other approaches to data fusion available in the literature include factor analysis (Cudeck,, 2000), statistical matching (Mitsuhiro and Hoshino,, 2020), Bayesian network inference (Triantafillou et al.,, 2010; Tsamardinos et al.,, 2012) and Gaussian Markov combinations (Massa and Riccomagno,, 2017). These methods are designed to link variables that are not simultaneously observed through covariates that exist for both variables of interest. This corresponds to the characteristics of in vivo data described above. However, the covariates in these approaches are continuous random variables, often assumed to be Gaussian, as is the case in Cudeck, (2000) and Massa and Riccomagno, (2017). The grouping variable available from in vivo experiments cannot be represented in continuous form, since categories such as control and sham make it impossible to assume continuity and normality. The Bayesian network approaches introduced by Triantafillou et al., (2010) and Tsamardinos et al., (2012), which aim to infer binary causal relationships between variables, are more suitable for large datasets with a high number of covariates. Current research in statistical matching addresses aspects such as not-at-random missingness (Mitsuhiro and Hoshino,, 2021) and high dimensionality (Mitsuhiro and Hoshino,, 2020). This approach is based on the idea of comparing distances between covariates from the datasets of interest, which cannot be done by taking the group variable as a covariate. It can be noted that the goal of the aforementioned examples in statistical matching is to group individuals prior to imputation, which is not necessary in our case since the groups are already known. Finally, ecological regression-based approaches are also employed, where the correlations of interest between two covariates relate to their means or percentages within groups. The primary difference with our study design is that we have access to the individual marginal observations of each variable. This allows us to conduct statistical inferences that ecological regression does not permit (Gelman et al.,, 2001).
In this work, we assume that there is a linear relationship between the continuous variables that are not simultaneously observed, and that the linear regression coefficients are the same for all sub-populations defined by the categorical variable. A similar context is treated in Evans et al., (2021), where a more general model is considered. The authors propose an approach that requires correctly specifying various relationships between variables (e.g., the distribution of the predictor variable conditional on what corresponds to the grouping variable in our case) in order to perform successful inference in the general case. Their approach is illustrated using survey data, which is the case when the latter can be expected to be successful due to large samples and/or prior knowledge. However, this is not the case for the experimental data considered in this work, which are characterized by small sample sizes and lack of prior knowledge about the underlying distributions. To address the specificities of the considered context, our approach is based on the assumption that there is a linear relationship conditional on the group, and that this relationship is the same for all groups, without making any assumptions about the distributions of the variables or requiring their specification.
We propose two estimators derived with the method of moments as well as an optimal transport solution using Wasserstein distance. Both approaches do not require any overlap between the two cohorts of the experiment and are based on weak assumptions ensuring model identifiability and the existence of moments. These estimators are shown to be consistent and asymptotically Gaussian under weak hypotheses. The asymptotic variance has no explicit expression, except in some special simple cases. For that reason a consistent stratified bootstrap approach is developed to construct confidence intervals for the estimated parameters. Not previously considered in the missing data literature, the bootstrap-based approach can be seen as the most important contribution of this paper, since the asymptotic solution is often infeasible in practice, whereas the bootstrap shows practical advantages in many realistic settings, especially in the case of a small sample size.
@@ 2 Identification approaches
We consider a real random variable $Y$ and a vector of $d$ real valued random regressors $\mathbf{X}=(X_{1},\ldots,X_{d})$, and suppose that the following linear regression holds:
$Y$ $=\beta_{0}+\sum_{j=1}^{d}\beta_{j}X_{j}+\epsilon.$ (1)
@@
In the absence of additional information and without a strong additional hypothesis, the parameters $(\beta_{0},\beta_{1},\ldots,\beta_{d})$ and the variance of the noise $\sigma^{2}_{\epsilon}$ cannot be identified. For example, if $X_{1}$ is centered and has a symmetric distribution, the coefficient $\beta_{1}$ can only be determined up to the sign change, since $\beta_{1}X_{1}$ and $\beta_{1}(-X_{1})$ have the same distribution.
To deal with this identification problem, we consider that we can perform different experiments in which the mean of $X$ is allowed to vary. To do this, we assume that there are $K$ groups (i.e. $K$ different experiments), defined by a categorical variable $G$ taking values in $\{1,\ldots,K\}$ observed simultaneously with $Y$ and $X$. This means that we now have access to $(\mathbf{X},G)$ and $(Y,G)$, but not to $(\mathbf{X},Y,G)$. We also assume that $\epsilon$ is independent of $G$.
Given $G=k$, for $k=1,\ldots,K$, we denote by $\mu_{Y}^{k}={\mathbb{E}}(Y|G=k)$ and $\mu_{X_{j}}^{k}={\mathbb{E}}(X_{j}|G=k)$, $j=1,\ldots,d$, the expected values within each group. We present two different approaches to identify the vector $\boldsymbol{\beta}=(\beta_{0},\ldots,\beta_{d})$ of unknown regression coefficients and the noise variance $\sigma_{\epsilon}^{2}$, taking into account the additional information provided by the discrete variable $G$.
@@ 2.1 Moment approach
The first simple approach is based on the first moments identification. Taking the conditional expectation, given $G=k$, in (1), we have for $k=1,\ldots,K$,
$\mu_{Y}^{k}$ $=\beta_{0}+\sum_{j=1}^{d}\beta_{j}\mu_{X_{j}}^{k},$ (2)
since the residual term $\epsilon$ is assumed to satisfy $\mathbb{E}(\epsilon|G=k)=0$ for $k=1,\ldots,K$.
We denote by $\boldsymbol{\mu}_{1,X}$ the $K\times(d+1)$ design matrix, with the $k$th row equal to $(1,\boldsymbol{\mu}_{X}^{k\top})$ with $\boldsymbol{\mu}_{X}^{k}=(\mu_{X_{1}}^{k},\cdots,\mu_{X_{d}}^{k})^{\top}$, and by $\boldsymbol{\mu}_{Y}$ the $K$ dimensional vector with elements $(\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$. The $K$ linear equations in (2) can be written in a matrix form: $\boldsymbol{\mu}_{Y}=\boldsymbol{\mu}_{1,X}\boldsymbol{\beta}$.
The following assumption guarantees the identifiability of the model parameters:
@@
Lemma 2.1
If the model (1) holds and the assumption $\mathbf{H}_{1}$ is fulfilled, $\boldsymbol{\beta}$ is uniquely identified in terms of the conditional first order moments of $\mathbf{X}$ and $Y$ given $G$,
@@ 2.2 Optimal transport approach
The second approach is based on optimal transport, in particular on the idea of estimating the linear transformation of the distribution of $\mathbf{X}$ that is the closest to that of $Y$ with respect to the Wasserstein distance (see Panaretos and Zemel, (2019) for a general introduction for statisticians). The optimal transport map T between Gaussian measures on $\mathbb{R}^{d}$ is linear, and the Wasserstein distance of order 2 between two Gaussian distributions $D_{1}$ and $D_{2}$, with $D_{1}\sim\mathcal{N}(\boldsymbol{\mu}_{1},\boldsymbol{\Gamma}_{1})$ and $D_{2}\sim\mathcal{N}(\boldsymbol{\mu}_{2},\boldsymbol{\Gamma}_{2})$, is equal to
@@
If the linear model (1) holds, and if we assume that, given $G=k$, $\mathbf{X}$ is a Gaussian random vector and $\epsilon$ is Gaussian, we have that $Y$ is also Gaussian given $G=k$, with expectation $\mu_{Y}^{k}=\beta_{0}+\sum_{j=1}^{d}\beta_{j}\mu_{X_{j}}^{k}$ and variance $\sigma_{\epsilon}^{2}+\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}$, where $\boldsymbol{\Gamma}_{X}^{k}$ is the variance matrix of $\mathbf{X}$ given $G=k$. Thus, the Wasserstein distance between $D_{\gamma}$, the distribution of $\gamma_{0}+\boldsymbol{\gamma}_{-0}^{\top}\mathbf{X}+\epsilon$, and $D_{Y}$, the distribution of $Y$, is equal to
@@
$\varphi(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\mu_{Y}^{k}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k})^{2}+\left(\sigma_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (3)
which evaluates the weighted Wasserstein distance between $Y$ and a linear combination of the $\mathbf{X}$ variables, contaminated by Gaussian noise, with variance $\sigma^{2}$. We state, without proof, the following lemma which ensures that the parameters $\boldsymbol{\beta}$ and $\sigma_{\epsilon}^{2}$ can be identified under general conditions:
Lemma 2.2
If the model (1) holds and the assumption $\mathbf{H}_{1}$ is fulfilled, $\varphi(\boldsymbol{\gamma},\sigma^{2})$ has its unique minimum at $\boldsymbol{\gamma}=\boldsymbol{\beta}$ and $\sigma^{2}=\sigma_{\epsilon}^{2}$.
@@ 3 Sampled data and estimators
We assume that the experiments are performed for $K\geq 2$ different groups, and that for each group $k$, for $k=1,\ldots,K$ we have two independent samples $(Y^{k}_{1},\ldots,Y^{k}_{n_{y}^{k}})$ and $(X^{k}_{j,1},\ldots,X^{k}_{j,n_{x}^{k}})_{j=1,\ldots,d}$, with sizes $n_{y}^{k}$ and $n_{x}^{k}$. For each unit $i=1,\ldots,n_{x}^{k}$ from group $k$, the vector of covariates is denoted by $\mathbf{X}_{i}^{k}=(X_{1,i},\ldots,X_{d,i})$. We also define $N_{x}=\sum_{k=1}^{K}n_{x}^{k}$ and $N_{y}=\sum_{k=1}^{K}n_{y}^{k}$, the total number of observations of the response $Y$ and the covariates $X_{1},\ldots,X_{d}$.
As shown in Lemma 2.1 and Lemma 2.2, the identification of the parameter $\boldsymbol{\beta}$ depends on the knowledge of the first two conditional moments, given $G$, of $Y$ and $\mathbf{X}$. For $k=1,\ldots,K$, we denote by $\widehat{\mu}_{Y}^{k}=\frac{1}{n_{y}^{k}}\sum_{i=1}^{n_{y}^{k}}Y^{k}_{i}$ and $\widehat{\mu}_{X_{j}}^{k}=\frac{1}{n_{x}^{k}}\sum_{i=1}^{n_{x}^{k}}X^{k}_{j,i}$ the empirical mean within each group $k$ and by $\widehat{\sigma}^{2}_{Y,k}=\frac{1}{n_{y}^{k}}\sum_{i=1}^{n_{y}^{k}}\left(Y_{i}^{k}\right)^{2}-\left(\widehat{\mu}_{Y}^{k}\right)^{2}$, and $\widehat{\boldsymbol{\Gamma}}^{k}_{X}=\frac{1}{n_{x}^{k}}\sum_{i=1}^{n_{x}^{k}}\mathbf{X}_{i}^{k}(\mathbf{X}_{i}^{k})^{\top}-\widehat{\boldsymbol{\mu}}_{X}^{k}(\widehat{\boldsymbol{\mu}}_{X}^{k})^{\top}$, the empirical variances, with $\widehat{\boldsymbol{\mu}}_{X}^{k}=(\widehat{\mu}_{X_{1}}^{k},\ldots,\widehat{\mu}_{X_{d}}^{k})$. We also define the total empirical mean and variance $\widehat{\mu}_{Y}=\frac{1}{N_{y}}\sum_{k=1}^{K}{n_{y}^{k}}\widehat{\mu}_{Y}^{k}$, $\widehat{\boldsymbol{\mu}}_{X}=\frac{1}{N_{x}}\sum_{k=1}^{K}{n_{x}^{k}}\widehat{\boldsymbol{\mu}}_{X}^{k}$, $\widehat{\sigma}^{2}_{Y}=\frac{1}{N_{y}}\sum_{k=1}^{K}\sum_{i=1}^{n_{y}^{k}}\left(Y_{i}^{k}\right)^{2}-\left(\widehat{\mu}_{Y}\right)^{2}$, $\widehat{\boldsymbol{\Gamma}}_{X}=\frac{1}{N_{x}}\sum_{k=1}^{K}\sum_{i=1}^{n_{x}^{k}}\mathbf{X}_{i}^{k}(\mathbf{X}_{i}^{k})^{\top}-\widehat{\boldsymbol{\mu}}_{X}\widehat{\boldsymbol{\mu}}_{X}^{\top}$. We denote by $\widehat{\boldsymbol{\mu}}_{1,X}$ the $K\times(d+1)$ matrix, with the first column consisting of ones, and the rest equal to $\widehat{\boldsymbol{\mu}}_{X}$.